OpenStudy (anonymous):

solve using differentiation please. a closed box, whose length is twice its width, is to have a surface 192 square inches. find the dimension of the box when the volume is maximum. answer is 4x8x5 1/3 in. solution please. :)

OpenStudy (anonymous):

Find a formula for the volume of the box. Use the surface area to express the unknown variable in terms of x. Differentiate the volume with respect to x. Set this to 0 to find the max value.

OpenStudy (anonymous):

value of x when f ' (x) = 0 = maximum value

OpenStudy (anonymous):

i still don't get the write answer. i have tried differentiating the volume with respect to x. still it's wrong.

OpenStudy (anonymous):

The total surface area of a box of width W, length L and height H is the surface area of the six rectangular sides, therefore\[2(WL+LH+WH)=192.\]Additionally, it is given that L=2W. Putting these together and doing a little rearranging gives \[2W^2 +3WH = 96\]and solving for H gives:\[ H = (96-2W^2)/3W\]The volume of the box is WLH, so substituting in our equations for H and L in terms of W:\[Volume = (192W-4W^3)/3 = 64W - (4/3)W^3\]Differentiate V with respect to W to get:\[dV/dW = 64 - 4W^2\]As noted above, the maximum volume is when dV/dW is set to zero, so:\[64-4W^2=0\]Solving for W gives W is 4, and hence L is 8 and H is 5.33 by substituting back into the equations derived above.

OpenStudy (anonymous):

you are so great! many thanks for the complete solution. :)

OpenStudy (anonymous):

No problem - the same method should work with any similar question about volume/surface area maximisation too :)